steam at 120 degree C is continuously passed through a 50cm long rubber tube of inner and outer radii 1cm and 1.2cm. the room temperature is 3 degree C.calculate the rate of heat flow through the walls of the tube.thermal conductivity of rubber= 0.15J/m-s-C

Grade:11

1 Answers

Chetan Mandayam Nayakar
12 years ago

T(r),1<r<1.2, dQ/dt=(k2pilr)(dT/dr),(2pilk)rdT/dr=constant(c),(2pilk)dT=dr/r,(2pilk)(T-276)=ln(r)c,c=(2pilk)*117/ln(1.2)=dQ/dt